OpenStudy (anonymous):

If the diagonal of a square is √30, what are the steps needed to find the lengths of the sides?

OpenStudy (anonymous):

divide by \(\sqrt2\) and get \(\sqrt{15}\)

OpenStudy (anonymous):

Ohhhh, right. Thank you!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

aliter assume each side to be a and use pythogoras theorem

OpenStudy (anonymous):

as a^2 +a^2 = (√30)^2 and further...

OpenStudy (mathstudent55):

 a^2 + a^2 = (sqrt(30))^2 2a^2 = 30 a^2 = 15 a = sqrt(15)

OpenStudy (anonymous):

Ok. That makes sense. Thank you! :)